for the first reaction process A+B-->product , the rate is first order with respect to A and second order with respect to B. if 1 mol each of A and B were introduced into a 1L besak, and the initial rate were 1×10^-2 mol/L /sec . calculate the rate when half the reactants have been turned into products

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
A +B = Product
Initial number of moles of A = 1 mol and For B = 1 mol
Initial volume = 1 litre
Injtial rate = 10^-2
Thus the rate equation can be written as-
Rate = k [a]¹ [b]²
Where k is the rate constant
k = 10^-2
Now when half of the reactants have been converted into product.
a = 0.5, b = 0.5
Thua the rate will be
Rate = k [0.5]¹ [0.5]²
= k/8
= 10^-2/8 = 125 * 10^-5